# Friedrich-Hirzebruch-Allee 5

### v1

@@ 7.1 Experiment Setup
Dataset: We evaluate MCDS-VSS on the Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] dataset, which contains 5,000 automotive video sequences, each containing 30 images of size 1024$\times$2048, recorded in 50 German cities, where only the 20th frame is annotated. This dataset is a good benchmark for our model, since it contains real-world dynamic scenes recorded from a moving vehicle. We augment the data using color jittering, mirroring, and random cropping.
@@
7.2 Comparison with Existing Methods
Table 2: Comparison of segmentation models on the Cityscapes validation set using different backbones. Best two results are highlighted in boldface and underlined, respectively.
@@ 7.2 Comparison with Existing Methods
[…]
Figure 3: Qualitative evaluation. a) Input frames, b) HRNetV2, c) MCDS-VSS (ours), d) Estimated scene depth, e) Estimated residual flow. We highlight areas of the segmentation masks where MCDS-VSS obtains visibly more accurate and consistent segmentations.
In Fig. 3, we show a qualitative result comparing MCDS-VSS with an HRNetV2 baseline. Our method achieves more accurate and temporally consistent segmentations compared to HRNet, correctly segmenting the traffic signs and reducing the amount of flickering between frames. Furthermore, we show MCDS-VSS interpretable intermediate representations, such as the depth of the scene and the residual flow, which encodes the movement of the pedestrians in the scene, as well as some corrections for other objects, such as the hood of the ego-vehicle. Further visualizations can be found in Appendix D.
@@ 7.3 Ablation Study
, we compare MCDS-VSS with different filter designs, including unstructured RNNs (ConvGRU) [Siam et al.(2017)Siam, Valipour, Jagersand, and Ray, Pfeuffer et al.(2019)Pfeuffer, Schulz, and Dietmayer], optical-flow based filters [Gadde et al.(2017)Gadde, Jampani, and Gehler] (flow-only), and a MCDS-VSS variant modeling only the static scene features (geom-only) [Wagner et al.(2018)Wagner, Fischer, Herman, and Behnke]. The results, reported in Table 3, show that MCDS-VSS, which decouples the modeling of static and dynamic scene features, achieves the best segmentation performance and temporal consistency.
Model Ablation: In Table 4 we measure the effect that our SSL training and MCDS-VSS filter have on the segmentation performance and temporal consistency. For two different segmentation models, i.e. DeepLabV3+ and HRNetV2, we compare the results after each training stage with those of the model trained for image segmentation only. First, we note that SSL training improves the temporal consistency without significantly compromising the segmentation performance. We argue that jointly learning SSL depth and motion with semantic segmentation allows the model to encode the input frames into more robust geometry-aware representations. Finally, the MCDS-VSS filter significantly improves the temporal consistency (>4.6% w.r.t base model), while almost matching the segmentation performance of the base DeepLabV3+ model, and even outperforming HRNetV2.
@@ 9 Conclusion
We proposed MCDS-VSS, a structured recurrent model for VSS, which learns in a self-supervised manner to estimate scene geometry and camera ego-motion. It also estimates the motion of external objects and leverages these representations to improve the temporal consistency of a semantic segmentation model without sacrificing segmentation performance. MCDS-VSS follows a prediction-fusion approach in which scene geometry and camera motion are first used to compensate for ego-motion, then residual flow is used to compensate the motion of dynamic objects, and finally the projected features are fused with the current observations. In our experiments, we showed that MCDS-VSS outperforms multiple VSS baselines on Cityscapes—achieving superior temporal consistency and parsing the scene into human-interpretable representations.
Acknowledgments
@@ B.1 Network Details
Residual Flow Decoder: The residual flow decoder is implemented as a modified lightweight version of RAFT [Teed and Deng(2020)]. To seamlessly integrate this module into our MCDS-VSS filter, we modify the implementation of raft_small provided by PyTorch [^2] https://pytorch.org/vision/main/models/raft.html by replacing the expensive feature and context encoders with a single convolutional block that directly processes the ego-warped $\mathbf{s}^{\textrm{ego}}_{t}$ and image features $\mathbf{h}_{t}$. We set the number of refinement iterations to 10.
@@
Table 7: Training stages and hyper-parameters.
@@
1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$ $2\cdot 10^{-4}$ 12 40h 3
@@ 3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$ $8\cdot 10^{-5}$ 8 8h 6
4 Temporal Integration $\mathcal{L}_{\textrm{Seg}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$ $8\cdot 10^{-5}$ 4 12h 6
B.2 Training and Inference
@@
Appendix C Additional Quantitative Results
@@
Appendix D Additional Qualitative Results
@@ D.1 Effect of Each MCDS-VSS Stage
In Fig. 4 we display the semantic segmentations obtained when decoding the scene features from different stages of our MCDS-VSS filter. We can observe how the segmentations after ego-motion compensation (Fig. 4 b) atone for the movement of the ego-vehicle, correctly representing static scene features such as buildings or the bicycle. However, the dynamics of moving objects (e.g. yellow car) are not addressed in this step, thus not compensating for such movement. This limitation is addressed in the object motion compensation step (Fig. 4 c). However, the disocclusions resulting from the moving car and inaccuracies in the residual flow estimation can lead to segmentation errors. Finally, fusing the projected scene state features with the current observations (Fig. 4 d) leads to a more accurate segmentation of the scene. In (Fig. 4 e) we visualize the update gate mask $\mathbf{u}$ employed for feature fusion. For visualization purposes, we take the mean across all channels and assign lighter colors to spatial locations where MCDS-VSS relies on the propagated state $\mathbf{s}^{\textrm{full}}_{t}$, whereas darker colors correspond to the current features $\mathbf{h}_{t}$. We observe that MCDS-VSS relies on the scene state to represent static areas such as buildings or the street, whereas it relies heavier on observations for accurately segmenting fast moving objects or thin structures (e.g. poles or street signs).
[…]
@@
[…]
@@
D.2 MCDS-VSS Qualitative Results
Despite being trained without any depth or motion labels, and solely with semantic segmentation annotations for a single frame per video, our model parses the scene into semantic, geometry and motion representations. In Figs. 58, we show on four validation sequences how MCDS-VSS obtains an accurate and temporally consistent segmentation of the scene, estimates the scene depth, and computes the residual motion flow, which encodes the movement of dynamic objects in the scene, as well as some minor motion corrections for other objects and scene features.
D.3 Point Clouds
Figs. 912 show examples of RGB and semantic point clouds rendered by backprojecting image values and semantic labels using the depth maps estimated by MCDS-VSS and known camera intrinsics. The high-quality depth maps computed by MCDS-VSS allow for an accurate 3D representation of the scene.
[…]

### v2

@@ 7.1 Experiment Setup
Dataset: We evaluate MCDS-VSS on the Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] dataset, which contains 5,000 automotive video sequences recorded in 50 German cities. Each sequence contains 30 images of size 1024$\times$2048, where only the 20th frame is annotated. This dataset is a good benchmark for our model, since it contains real-world dynamic scenes recorded from a moving vehicle. We augment the data using color jittering, mirroring and random cropping.
@@
Table 2: Comparison of image and video segmentation models on the Cityscapes validation set using small (left) and larger (right) backbones. We evaluate the segmentation accuracy (mIoU) and temporal consistency (TC) of the models. Best two results are highlighted in boldface and underlined, respectively.
@@
7.2 Comparison with Existing Methods
Figure 3: Qualitative evaluation on a validation sequence of five frames. a) Input frames, b) HRNetV2, c) MCDS-VSS (ours), d) Estimated scene depth, e) Estimated residual flow. We highlight areas of the segmentation masks where MCDS-VSS obtains visibly more accurate and temporally consistent segmentations, such as the traffic signs or the bus, which HRNetV2 mislabels as truck.
In Figure 3, we show a qualitative result comparing MCDS-VSS with the HRNetV2 baseline on a validation sequence of five frames. Whereas the baseline mislabels the bus as a truck and outputs inconsistent segmentation labels on certain regions such as the traffic signs, our method achieves more accurate and temporally consistent segmentations, predicting more stable semantic labels across video frames. Furthermore, we show MCDS-VSS interpretable intermediate representations, such as the estimated scene depth and residual optical flow, which encodes the movement of the vehicles in the scene, as well as corrections for the hood of the ego-vehicle. Further visualizations can be found in Appendix D.
@@ 7.3 Ablation Study
, we compare MCDS-VSS with different filter designs, including unstructured RNNs (ConvGRU) [Siam et al.(2017)Siam, Valipour, Jagersand, and Ray, Pfeuffer et al.(2019)Pfeuffer, Schulz, and Dietmayer], optical-flow based filters [Gadde et al.(2017)Gadde, Jampani, and Gehler] (flow-only), and a MCDS-VSS variant modeling only the static scene features (geom-only) [Wagner et al.(2018)Wagner, Fischer, Herman, and Behnke]. The results, reported in Table 3, show that filter designs that project scene features using geometry and motion representations outperform the ConvGRU, which learns to model the video dynamics solely from data. Furthermore, MCDS-VSS, which decouples the modeling of static and dynamic scene features, achieves the best segmentation performance and temporal consistency among the compared filter designs.
Model Ablation: In Table 4 we measure the effect that our joint training procedure of semantic segmentation and SSL depth and ego-motion, as well as the MCDS-VSS filter have on the segmentation performance and temporal consistency. For two different segmentation models, i.e. DeepLabV3+ with a ResNet18 backbone and HRNetV2, we compare the results after each training stage with those of the model trained for image segmentation only. First, we note that jointly learning semantic segmentation with SSL depth and ego-motion estimation improves the temporal consistency without significantly compromising the segmentation performance. We argue that the joint training procedure allows the model to encode the input frames into more robust geometry-aware representations. Finally, the MCDS-VSS filter significantly improves the temporal consistency (>4.6% w.r.t base model), while almost matching the segmentation performance of the base DeepLabV3+ model, and even outperforming HRNetV2.
@@ 9 Conclusion
We proposed MCDS-VSS, a structured recurrent model for VSS, which learns in a self-supervised manner to estimate scene geometry and camera ego-motion. It also estimates the motion of external objects and leverages these representations to improve the temporal consistency of a semantic segmentation model without sacrificing segmentation performance. MCDS-VSS follows a prediction-fusion approach in which scene geometry and camera motion are first used to compensate for ego-motion, then residual flow is used to compensate the motion of dynamic objects, and finally the projected features are fused with the current observations in order to obtain a temporally consistent representation of the scene. In our experiments, we showed that MCDS-VSS outperforms multiple VSS baselines on Cityscapes—achieving superior segmentation temporal consistency and parsing the scene into human-interpretable representations, such as depth, ego-motion and object flow.
Acknowledgment
@@ B.1 Network Details
Residual Flow Decoder: The residual flow decoder is implemented as a modified lightweight version of RAFT [Teed and Deng(2020)]. To seamlessly integrate this module into our MCDS-VSS filter, we modify the implementation of raft_small provided by PyTorch [^2] https://pytorch.org/vision/main/models/raft.html by replacing the expensive feature and context encoders with a single convolutional block that directly processes the ego-warped $\mathbf{s}^{\textrm{ego}}_{t}$ and image features $\mathbf{h}_{t}$. We set the number of refinement iterations to 12.
@@
B.2 Training and Inference
Table 7: MCDS-VSS training stages and hyper-parameters.
@@ B.2 Training and Inference
1 Segmentation & SSL Geometry $\mathcal{L}_{\textrm{Segm}}+\lambda_{\textrm{D}}\cdot\mathcal{L}_{\textrm{Depth}}$ $2\cdot 10^{-4}$ 12 40h 3
@@ 3 Ego-Motion Filter $\mathcal{L}_{\textrm{Ego}}$ $8\cdot 10^{-5}$ 8 8h 6
4 Temporal Integration $\mathcal{L}_{\textrm{Segm}}+\lambda_{\textrm{TC}}\cdot\mathcal{L}_{\textrm{TC}}$ $8\cdot 10^{-5}$ 4 12h 6
@@
Appendix C Quantitative Results
@@
Appendix D Qualitative Results
@@ D.1 Effect of Each MCDS-VSS Stage
In Figure 4 we display the semantic segmentations obtained when decoding the scene features from different stages of our MCDS-VSS filter. We can observe how the segmentations after ego-motion compensation (Figure 4 b) atone for the movement of the ego-vehicle, correctly representing static scene features such as buildings or the bicycle. However, the dynamics of moving objects (e.g. yellow car) are not addressed in this step, thus not compensating for such movement. This limitation is addressed in the object motion compensation step (Figure 4 c). However, the disocclusions resulting from the moving car and inaccuracies in the residual flow estimation can lead to segmentation errors. Finally, fusing the projected scene state features with the current observations (Figure 4 d) leads to a more accurate segmentation of the scene. In (Figure 4 e) we visualize the update gate mask $\mathbf{u}$ employed for feature fusion. For visualization purposes, we take the mean across all channels and assign lighter colors to spatial locations where MCDS-VSS relies on the propagated state $\mathbf{s}^{\textrm{full}}_{t}$, whereas darker colors correspond to the current features $\mathbf{h}_{t}$. We observe that MCDS-VSS relies on the scene state to represent static areas such as buildings or the street, whereas it relies heavier on observations for accurately segmenting disoccluded areas of the image, fast moving objects or thin structures (e.g. poles or street signs).
@@
D.2 MCDS-VSS Qualitative Results
In Figure 5, we show a qualitative result comparing MCDS-VSS with the HRNetV2 baseline. Our method achieves more accurate and temporally consistent segmentations compared to HRNet, correctly segmenting the traffic signs and reducing the amount of flickering between frames.
In Figures 69, we show on four validation sequences how MCDS-VSS obtains an accurate and temporally consistent segmentation of the scene, estimates the scene depth, and computes the residual motion flow, which encodes the movement of dynamic objects in the scene, as well as some minor motion corrections for other objects and scene features.
D.3 Cross-Dataset Evaluation
We qualitatively evaluate the robustness of MCDS-VSS by performing a cross-dataset validation in which a DeepLabV3+ baseline and our MCDS-VSS model trained on Cityscapes are qualitatively evaluated without retraining on sequences from the KITTI [Geiger et al.(2012)Geiger, Lenz, and Urtasun] dataset. Figures 10 and 11 illustrate the semantic segmentation predictions of both models, as well as the MCDS-VSS depth estimates on two validation sequences of the KITTI dataset. Due to differences with respect to the training data in the camera model and calibration, as well as different image resolution and aspect ratio, the segmentation performance of both models on the KITTI dataset is severely degraded with respect to Cityscapes. However, MCDS-VSS achieves a more accurate and temporally consistent video segmentation, thus verifying that incorporating geometry and motion inductive biases from the moving camera dynamic scene domain into the VSS model design leads to more robust representations and segmentation results.
D.4 Point Clouds
Figures 1215 show examples of RGB and semantic point clouds rendered by backprojecting image values and semantic labels using the depth maps estimated by MCDS-VSS and known camera intrinsics. The high-quality depth maps computed by MCDS-VSS allow for an accurate 3D representation of the scene.
Figure 5: Qualitative evaluation. a) Input frames, b) HRNetV2, c) MCDS-VSS , d) Estimated scene depth, e) Estimated residual flow. We highlight areas of the segmentation masks where MCDS-VSS obtains visibly more accurate and temporally consistent segmentations.
@@ D.4 Point Clouds
Figure 10: Cross-dataset qualitative evaluation of models trained on Cityscapes and evaluated on KITTI. a) Input frames, b) DeepLabV3+ baseline, c) MCDS-VSS (ours), d) estimated scene depth.
Figure 11: Cross-dataset qualitative evaluation of models trained on Cityscapes and evaluated on KITTI. a) Input frames, b) DeepLabV3+ baseline, c) MCDS-VSS (ours), d) estimated scene depth.